### v3

# Distributional Scaling of Emergent Capabilities

Abstract
This paper explores the nature of sudden breakthroughs in language model performance at scale, which stand in contrast to smooth improvements governed by scaling laws. While advocates of “emergence" view breakthroughs as unlocked capabilities, others attribute them to thresholding effects on noncontinuous metrics. We propose that breakthroughs are instead driven by continuous changes in the probability distribution of training outcomes when performance is bimodally distributed across random seeds. In synthetic length generalization tasks, we show that different random seeds can produce either highly linear or emergent scaling trends. We reveal that sharp breakthroughs in metrics are produced by underlying continuous changes in their distribution across seeds. Furthermore, we provide a case study of inverse scaling and show that even as the probability of a successful run declines, the average performance of a successful run continues to increase monotonically. We validate our distributional scaling framework on realistic settings by measuring MMLU performance in LM populations. These insights emphasize the role of random variation in the effect of scale on LM capabilities.
1 Introduction
On most benchmarks, language model (LM) performance is determined by a scaling law [12, 35, 21] that varies smoothly with parameter size and overall training compute. There are, however, a number of celebrated exceptions in which performance abruptly improves on specific benchmarks [41].
These sudden breakthroughs fuel one of the most heated debates in modern machine learning. On one side, advocates of emergence claim that performance abruptly improves at particular scales because those scales allow the LM to acquire specific concepts that permit out-of-distribution generalization [45]. On the other side, skeptics argue that these sudden improvements are a mirage driven by thresholding effects and alleviated by more appropriate continuous metrics—though a few breakthrough capabilities remain stubbornly emergent [36]. Here, we argue that such discontinuities are driven by continuous changes in the probability of a breakthrough at each scale.
We posit that a breakthrough capability is often distinguished not by deterministic responses to scale, but by multimodal random variation—in other words, independent training runs form multiple clusters in their performance metrics. Such effects are currently undetected because scaling laws are usually determined by measuring a single training run at each scale, and resources are rarely committed to correct for random variation by reusing the same hyperparameter settings with different random seeds. Although random variation may be benign when model performance is measured in-distribution [18], previous work suggests that out-of-distribution performance may vary widely across training runs [49, 51], even at larger scales [26]. Since training numerous seeds becomes prohibitively expensive at large scales, we first study length generalization in small models trained on synthetic tasks. In similar settings, existing publications consider only a few training runs and report their summary statistics. By contrast, we characterize their full distribution, demonstrating that these breakthrough tasks exhibit multimodal variation. Our findings can be summarized as follows:
Figure 1: Different random seeds produce different scaling trends. Scaling trends can be emergent or linear for different seeds, even if all models train on the same data with the same hyperparameters. On the counting task, we show trends for random seeds with the highest breakthroughness (seed 93; top left) and linearity (seed 205; top right). We mark parameter counts immediately before and after seed 93’s emergence respectively as (a) and (b). Histograms illustrate the bimodal distribution of performance across all random seeds at scales (a) and (b), marking the positions of seeds 93 and 205. Breakthroughs occur when consecutive points represent different clusters; linear trends occur when each point is sampled from the same gradually shifting cluster.
  • Breakthrough scaling results from bimodal performance distributions. In language modeling and synthetic length generalization case studies, we demonstrate that some random seeds produce linear scale trends while others are emergent. This variation is caused by the bimodal distribution of each skill across random seeds, a property that materializes around breakthrough thresholds in model size. At these scales, “emergence” is a stochastic property. We confirm these capabilities remain bimodally distributed under continuous metrics.
  • While a scale curve exhibits sudden, discontinuous improvement in a skill, the probability of learning that skill may be responding continuously to scale. Modeling the bimodal distribution as a mix of failure and success distributions, we illustrate that average improvements can come from changes in the probability of success or in the mean performance of a successful model (Section 3). Measuring the Wassterstein distances between distributions at each scale (Section 3.3), we show that bimodality appears abruptly at some minimum capacity scale, but these sudden changes do not necessarily align with emergent trends for individual model runs.
  • Competing solutions can lead to either monotonic or U-shaped trends in emergence likelihood. We discuss a synthetic task that exhibits periods of inverse scaling in average improvement (Section 3.4), a phenomenon also observed in LMs [28]. We note that the mean performance of successful runs remain monotonic during ranges of inverse scaling, suggesting that U-shaped curves in performance distributions are an artifact of success probability.
  • In realistic settings, multimodal performance clusters emerge based on data composition and model scale. In Section 4, we extend our analysis to the task of multiple choice question answering by continually pretraining LMs. Our findings confirm that random variation in LMs is also bimodal for established emergent capabilities near their breakthrough threshold.
(a) Fixing network depth to 6 layers.
(b) Fixing network width to 8 heads per layer.
Figure 2: Random variation in length generalization (addition task). Histograms of exact match accuracy on length $40$ sequences when independently scaling 2(a) width and 2(b) depth.
2 Experiments
2.1 Synthetic Tasks
We measure models’ length generalization [10, 20, 23, 15], one of many compositional properties that lead to emergence [41, 25, 3]. Experimental setup is described below, with details in Appendix A.
Architecture: In our synthetic experiments, we train decoder-only Transformer models from scratch using rotary position embeddings (RoPE) [42]. To observe the random performance distribution at each scale, we train our models from hundreds of seeds. We choose model sizes by separately increasing the width and depth hyperparameters, fixing the other hyperparameter. We scale width by increasing the number of 64-parameter heads per layer.
Task: We consider two algorithmic tasks previously studied in Zhou et al. [49]: counting and reverse order addition. Performance on these task ranges widely across random seeds. At each scale, we train models from 250 seeds for the count task and 200 seeds for reverse order addition.
  • Count: Given two numbers in increasing order, the model is trained to generate a sequence which counts consecutively from the first number to the second number. Examples are given in the form "5, 9 >, 5, 6, 7, 8, 9", while limiting the length of the counting sequence during training. Zhou et al. [49] showed that models trained to count can generalize to more than twice this training length; however, we will reveal a more nuanced view of length generalization based on its distribution across independent model runs.
  • Reverse Order Addition with Index Hints: Zhou et al. [51] showed that Transformers can generalize 10-15 digits past training length for an addition task, if provided with index hints and allowed to generate the answer backwards. Examples take the form "a0, 3, a1, 4, +, a0, 2, a1, 8, >, a1, 2, a0, 6", where index hints are a consecutive sequence of the max number of digits sampled randomly from 0 to the max evaluation length. In this work, we refer to this task as the ‘addition task’ despite these modifications.
Dataset: During training, we sample sequences i.i.d from the train set and invoke in-context learning by adding examples to the context, following previous work [17, 51]. The length of examples are sampled uniformly from 1 to the maximum training length, set to 30 for count and 35 for addition. Length generalization is then tested at length 60 for count and 40 for addition.
2.2 Natural Language Tasks
We extend our investigation to LMs by testing multiple choice question answering, an emergent natural language understanding task [41, 40]. (A second task, CommonSenseQA, is in Appendix E.) We test whether the performance clusters observed in emergent synthetic tasks also manifest in LMs. Specifically, we hypothesize that emergent scaling curves in LMs express underlying bimodal (or multimodal) performance distributions.
Our procedure can be regarded as continued pretraining, as we mimic the model’s original pretraining objective with a dataset of English language sequences. We hypothesize that at the emergence point for this task, performance distributions will also be bimodal across random seeds. During pretraining, a large diverse corpus encourages models to acquire various capabilities. While sufficiently large models may learn all such capabilities, smaller models have limited capacity, requiring these capabilities to compete [29]. This competition, influenced by initialization and data order, leads to varying outcomes across random seeds, forming performance clusters for specific benchmarks.
Task: We consider finetuning LMs for the MMLU benchmark [11], a multiple-choice question-answering task. Achieving strong performance on MMLU requires two key abilities: (1) natural language reasoning with domain-specific knowledge and (2) producing correct answers in the required format. It is this latter ability which leads to emergent trends [41, 13].
Model: We use the Qwen2.5 family of base models [47]. To introduce randomness during continued pretraining, we reinitialize the final attention layer and the subsequent LM head. We perform full-parameter continued pretraining on these reinitialized models.
Data: To ensure that the multiple choice formatting circuits compete with other tasks during continued pretraining, we create a diverse dataset by mixing the English news subset of C4 [34] with the official MMLU auxiliary training data. The C4 news data reinforces general language modeling ability, while the MMLU data focuses on multiple-choice question answering. We vary the proportion of MMLU in this mixture to control the target task’s data size in addition to model size, as both are common control variables in the scaling laws literature.
Training: We continue pretraining the Qwen2.5-0.5B and Qwen2.5-1.5B models on our C4-MMLU mixes, training 20 reinitializations on each data mixture ratio. We train Qwen2.5-0.5B models for 2 epochs and Qwen2.5-1.5B models for 5 epochs to ensure convergence on the MMLU validation set. We use a learning rate of 1e-5 with linear decay scheduler.
3 Synthetic Experiment Results
Existing studies on performance breakthroughs across scales typically report results for a single model or, at most, the average of a few runs. The literature suggests that these emergent behaviors are unlocked at certain model scales [45], implying that scaling curves for different model runs would generally perform similarly to within a margin of noise. Contrary to this belief, we demonstrate that breakthroughs in compositional tasks are the result of a stochastic distribution which changes gradually even as individual scaling curves jump abruptly. We will introduce these concepts and our framework through a case study of the addition task from Section 2.1.
3.1 Emergence is a Sign of Bimodal Variation
First, we report benchmark performance across scales on our synthetic tasks by emulating the conventional approach in reporting results for single seeds. Following Srivastava et al. [41], we take the vector of model performances for the count task given a fixed seed at test length 60 across different scales and calculate their breakthroughness and linearity metrics. As defined in Appendix B, the former metric measures emergence, whereas the latter measures a smooth response to scale. We plot the performance across scale for the seeds with the highest breakthrough and highest linearity in Figure 1, fixing the random initialization and data order seed across scales. [^1] Although it is standard practice to fix the random seed when reporting LM benchmark performance across scales, the initializations produced by a single seed have no meaningful relation across different scales. As seen in Figure 1 and Appendix Figure 8, we can easily find fixed seeds that show varying levels of emergence and linearity, due to random variation in breakthroughs.
This variation is explained by the bimodality of model performance distributions when varying seed. The histograms shown in Figure 2 illustrate that, for a population of models independently trained on the addition task, performance clusters into high and low component modes at many parameter sizes. In other words, these scales exhibit distinctly bimodal distributions in length generalization capabilities. This variability is precisely what causes some model runs to appear as breakthroughs while others follow a more linear progression. Specifically, when different runs are clustered into very high or very low performance due to their bimodality, a model might exhibit linear scaling if sampled from the same cluster as the previous scale or emergent scaling when switching from the low cluster to the high cluster. These differences ultimately lead to high variability in the timing and degree of emergence.
3.2 Sudden Jumps From Gradual Distribution Shifts
When reporting metrics from only one seed or the mean of a few seeds, the outcome is likely to be close to the mode performance of the underlying model population. As shown in Figure 3 (top left), the mode of the performance distribution shows a massive spike in improvement mirroring an emergent benchmark’s scaling trend. However, we claim that this discontinuous improvement is only an artifact of sampling the mode performance, although this artifact which will also produce discontinuity in many single-seed scaling curves. Underlying this discontinuous performance jump are continuous changes in other distributional statistics. In Figure 3 (top right), the mean exhibits a smoother trend in accuracy. Mode and mean diverge thus because the underlying distribution is bimodal, expressing a mixture of “successful” and “failing” runs.
Treating the distribution as a mixture of successes and failures, we can separately analyze the probability of a successful run and the performance distribution of successful runs. Both of these properties are changing continuously and gradually when the mode increases abruptly. If we restrict our analysis to the runs achieving nontrivial 20% accuracy, we see that the probability (Figure 3 (bottom left) and mean (bottom right) of such “successful” runs both exhibit continuous improvement, with the exception of increasing from depth 2 to 3 (discussed further in Section 3.3). Even at the mode breakthrough,these underlying distributional properties are only changing gradually. We conclude that when tasks exhibit bimodal distributions across random seeds, there are statistics exhibiting continuous improvements underlying the seemingly abrupt improvements across scale.
Figure 3: Summary statistics for length generalization (addition task). Exact match statistics for 200 models trained at length 35 and tested at length 40. We track overall 3 mode and 3 mean. Because the EM accuracy distribution is bimodal, the mode exhibits a sharp increase even as the mean evolves continuously. Defining success as >20% EM accuracy, we also note the continuous change in 3 the probability of success and in 3 the mean of successful runs . Means feature 95% confidence intervals over 1000 bootstrapped samples.
(a) Addition task. The sharpest change in distance occurs early. During the mode breakthrough, the distributions are still shifting gradually.
(b) Count task. The distribution changes slowly between intermediate model scales, but changes suddenly at the nadir of the U-shaped curve in Figure 5.
Figure 4: Changes in random variation. Wasserstein-L2 distance of each scale’s performance distribution relative to the largest scale, scaling depth and width independently. We mark the emergence of bimodality at the last scale before multiple peaks appear. We mark the mode breakthrough at the last scale before successful length generalization becomes marginally more likely than failure.
3.3 Bimodality Emerges Abruptly
The previous section showed that bimodal performance distributions produce emergent scaling curves for individual seeds. We next analyze how the distribution itself evolves across scales. As shown in Figure 2, the performance distribution starts as unimodal at the smallest scale, where no models can length generalize. Larger scales yield a bimodal distribution where most of the probability mass is ultimately placed on successful length generalizing runs. A priori, there could be a smooth evolution between these two distributions in which probability mass from failing runs gradually shifts towards higher performance metrics, eventually splitting into a clear separate cluster. We find instead that the shift from unimodal to bimodal is abrupt and instantly polarized into low and high clusters. In Appendix Figure 10 we show that models fail to generalize when fixing depth to be one layer even when scaling width (despite having perfect in-distribution accuracy).
To track the evolution of the performance distribution across scales, we plot the Wasserstein-L2 distance of each distribution relative to that of the final, largest model scale when separately fixing width and depth. In Figure 4(a) we see that there is a sharp decrease in the W2 distance for scaling with a fixed width, quantifying the sudden appearance of highly successful runs when model depth reaches 3 layers or model width reaches 2 heads. These sudden changes identify the moment when a new capability is unlocked, as the distribution transitions abruptly from unimodal on one extreme to bimodal at both extremes. We posit that this transition marks the minimum capacity required to learn the task. We also mark the point in each trend where the mode in Figure 3 (top left) increases sharply. We posit that it may be misleading to draw conclusions about minimal model capacity on a specific task using single runs at each scale, whereas distributional metrics correctly predict the sudden appearance of probability mass placed on successful runs.
3.4 Count: An Instance of U-shaped Scaling
We next consider the counting task (Section 2.1), which also exhibits bimodally-distributed performance (see Appendix Figure 9) but ultimately yields a very different scaling effect: a U-shaped curve. Figure 5 reveals this peculiar phenomenon in the mean accuracy scaling curve, when holding depth fixed. This curve is not simply a result of the summary statistic chosen, as it is mirrored by the evolution of the distributions as a whole according to their W2 distance in Figure 4(b).
U-shaped scaling has been observed in LMs, but its causes are not currently well-understood. When the Inverse Scaling Prize [28] solicited tasks which exhibit inverse scaling trends—performance decreases at scale—for large models, Wei et al. [46] revealed that the majority of awarded tasks actually exhibit U-shaped scaling after considering even larger models. Treating the counting task as a concrete instance of U-shaped scaling at small model scales, we find that this unusual trend is still underscored by monotonic continuous changes in the performance distribution. Indeed, Figure 5 (bottom right) shows that although the trend in the mean across all runs is U-shaped curve, the mean of the “successful” runs—those achieving at least 50% accuracy—still improves monotonically when increasing width. The observation of inverse scaling is, instead, due to changes in the probability of success (bottom left). Even when inverse scaling is in effect across a performance distribution, the performance of successful runs may exhibit more conventional responses to scale.
Figure 5: Summary statistics for length generalization (count task). Models tested at length 60, scaling width (with fixed depth 4) and depth (with fixed width 4) independently. The trend for fixed depth (scaling width) exhibits a U-shaped curve for the mean across all runs 5; however, the mean of successful runs (defined as >50% EM accuracy) 5 still exhibits continuous improvement.
3.5 Is Bimodality a Mirage?
Metrics with hard thresholds and discontinuities can artificially induce breakthrough behavior [36]; conversely, continuous metrics can make apparently emergent curves into smooth curves [41]. We must be particularly cautious about claiming emergence when requiring model outputs to exactly match a target string, as we do for both synthetic tasks. Are our case studies artifacts of thresholding effects? In other words, do their bimodal distributions become unimodal under continuous metrics? To assess the role of the exact match metric in creating apparent bimodality, we instead consider a continuous equivalent to the exact match metric: the minimum probability assigned to any individual token. If the model fails to assign its highest probability to the correct token for any position in the sequence, it receives a score of zero under the exact match metric, even if the correct token was still highly ranked. Likewise, if the model successfully assigns maximum probability to the correct token, exact match considers it to be accurate whether that probability was high or low. We now consider a continuous version of EM accuracy that reports the worst mistake made by the model. This continuous error score for model $f$ on dataset $X$ is based on the per-token loss $L(f(x_{0\ldots i-1}),x_{i})$:
$\textrm{error}(f,X)=\frac{1}{|X|}\sum_{x\in X}\max_{i<|x|}L(f(x_{0\ldots i-1}),x_{i})$ (1)
In Figure 6, we plot this continuous metric for addition models on the length generalization test dataset. The distribution of this metric across random seeds is often still clustered. We therefore confirm that emergent capabilities exhibit bimodal performance distributions even when using a continuous performance metric; their bimodality is not due to thresholding alone. Further discussion around continuous metric plots for count and histograms for probability-based metrics are in Appendix D.
(a) Fixing network depth to 6 layers while scaling up width.
(b) Fixing network width to 8 heads per layer while scaling up depth.
Figure 6: Random variation in continuous length generalization error (addition task). Kernel Density Estimation (KDE) of our loss-based error metric (Equation 1) across model runs. At scales where the EM accuracy distribution is bimodal, the distribution remains bimodal even when using a continuous metric.
(a) Qwen2.5-0.5B trained on different mixes.
(b) Qwen2.5 models trained on 10% MMLU data.
Figure 7: Performance on MMLU. All models have their last layer reinitialized randomly and then are trained on a mix of C4 news and MMLU, with varying proportions of MMLU. Left: With insufficient MMLU training data, each model fails to follow the multiple-choice format, resulting in trivial performance. As more MMLU data is included, different seeds lead to bimodally distributed performance. With enough MMLU data, all models score above the random baseline. Right: For a fixed data mix, the smaller model has bimodally distributed performance. In contrast, the larger model consistently relearns to answer multiple-choice questions, though not at its original base performance.
4 Natural Language Results
After connecting emergence with bimodality in small synthetic settings, we turn to LMs. We focus on the MMLU dataset, where high performance emerges after the LM learns the multiple choice format [41, 13]. To avoid the expense of repeatedly training large-scale models from scratch, we simulate independent runs by reinitializing the upper layer of pretrained LMs before continuing to train them. Because Qwen base models perform well for their scale (39.5% for Qwen2.5-0.5B), these experiments effectively require models to re-discover MMLU capabilities after their removal. Our continued training dataset, a mix of C4 news data and the MMLU auxiliary training data, is insufficient to recover the full MMLU capability of the the base model.
4.1 Emergence Across Data Compositions
We first examine how data composition influences the emergence of bimodal behavior in Qwen2.5-0.5B. Specifically, as the training data contains more examples of a target breakthrough task, it changes the task’s bimodal performance distribution in ways that mirror the effect of scale.
In Figure 7 (left), when trained on a data mix containing only 5% MMLU examples and 95% random sequences from C4 [34], most models (out of 20 seeds) achieve near 0% performance on the MMLU test set. According to the existing literature, this failure stems from a failure to process the multiple-choice format [13]. As the proportion of MMLU samples increases, bimodality emerges when MMLU composes 10% of the training data, where models form two distinct performance clusters. In one cluster, continually pretrained models still fail to follow the correct format, while in the other, models achieve performance at or above the random guess baseline of 25%. This second cluster contains models that consistently respond with valid multiple choice options; those that outperform the random baseline have even learned to compose the multiple choice format with their world knowledge when selecting an answer. Finally, with sufficient data ($>20\%$ MMLU), models across all 20 seeds consistently outperform the 25% random baseline.
4.2 Emergence Across Model Scales
Next, we show that, as in the synthetic setting, LM model size affects the formation of performance clusters. Fixing the dataset to a mixture with 10% MMLU auxiliary training samples, we continually pretrain Qwen2.5-1.5B on the same 20 seeds and report MMLU test accuracy in Figure 7 (right). While models built on the smaller Qwen2.5-0.5B form two distinct performance clusters, those built on the larger Qwen2.5-1.5B consistently outperform the random baseline. We conclude that larger scale models can reliably acquire MMLU capability when trained on the same dataset that produces highly bimodal variation at smaller scales. This conclusion recalls the literature on scaling laws, which argues that smaller models require more training examples to match the performance of larger models [35, 21]. In Appendix E.1, we show similar results in the Commonsense QA [43] task.
5 Related Works
Random variation: Model performance can be sensitive to stochastic aspects of the training process like random initialization and training data order. Prior studies have documented performance differences across various stress test sets [4, 30]. Such differences persist throughout training, not just at the final checkpoint [50]. Dodge et al. [6] compared the impacts of weight initialization and data ordering, concluding that both contribute equally to variation in performance. Additionally, Sellam et al. [38] demonstrated that LMs independently pre-trained from scratch can differ in downstream task performance and social biases. Existing work has also found model runs can cluster in out-of-distribution behavior [19] and in training dynamics [33, 14], hinting at multimodal variation. Our work connects these random clustering effects to the phenomenon of emergence at scale.
Length generalization and compositionality with transformers: Transformer models cannot always generalize to longer sequences than they are trained on [1, 5, 9, 15, 37, 48]. Various proposals have, consequently, aimed to improve length generalization. Some methods focus on alternative positional encodings [39, 32, 42, 22, 17]. Others modify the dataset format by adding scratchpad or Chain-of-Thought [1] or by incorporating padding and index hints [16, 49]. Regarding random variation, Zhou et al. [49] and Zhou et al. [51] provide evidence of variability in length generalization across random seeds, which we further investigate and analyze across a range of model scales.
Emergent abilities of LMs: In large language models, emergent abilities are behaviors that arise unexpectedly as models are scaled up in size or trained on larger datasets [12, 35, 2, 21]. These abilities are characterized by unpredictable and abrupt performance improvements on specific benchmarks at certain scales [45, 8, 41]. Understanding the conditions and mechanisms underlying emergence is a key area of research. Although recent studies suggest that some breakthroughs may stem from the choice of evaluation metrics rather than fundamental changes in model behavior with increased scale [41, 36], some breakthrough capabilities remain emergent. Snell et al. [40] found that some scales exhibit earlier emergence if finetuned explicitly on an emergent task, suggesting that smaller models may have the capacity for that task but are limited by its scarcity in the training corpus. Similarly, we show that emergent capabilities can arise from multimodal random variation using synthetic length generalization tasks as a case study.
Depth versus Width Scaling: Although scaling laws offer a smooth extrapolation for model performance, downstream performance can vary depending on architecture shape and not just model size [44]. For compositional tasks, deeper models often generalize better up to a certain point, but for a fixed compute budget, it may be more advantageous to train a shallower, wider model [31]. Various works have proposed explanations for the role of width versus depth in scaling behavior; for instance, Edelman et al. [7] show that increasing network width offers more ‘parallel queries’ over randomized subnetworks which learn sparse features more efficiently. Levine et al. [24] use a border rank argument to establish a width-dependent depth threshold, beyond which additional depth yields diminishing returns. In this work, we specifically investigate how independently scaling width and depth influences the random variation distribution in compositional tasks.
6 Discussion and Conclusions
Our work explores the evolution of random variation in model performance across scales, bringing a more nuanced perspective on emergent capabilities than the conventional approach of plotting a single model run per scale. Previous work has already documented the variability of length generalization across random seeds [51, 49]; in general, out-of-distribution behavior like compositional rules [27] or associative biases [38] often exhibit extreme variation compared to in-distribution performance. We attribute a capability’s emergence to its underlying bimodal random distribution, which is also documented in text classifier performance metrics [19] and even in the timing of generalization breakthroughs during training [14]. The mode of these performance distributions displays a sharp improvement at a certain model scale, similar to emergent capabilities in the literature, but we attribute these sudden jumps to gradual improvements in the random distribution. Furthermore, we show that bimodality can emerge before the mode—or most seeds—exhibit a breakthrough (Section 3.3); the transition from the initial unimodal distribution to a bimodal one is sudden.
In the wider scaling laws literature, which often focuses on aggregated loss, parameters added by increasing depth and width are often treated interchangeably in their contribution to total compute. By separately considering depth and width, our results suggest that emergent phenomena may be sensitive to tradeoffs between architectural hyperparameters. In particular, we document a regime in which increasing width damages model performance but increasing depth improves model performance. In a research community increasingly concerned with emergent capabilities, our findings should inspire further study of how these tasks respond to architectural hyperparameter tradeoffs.
Acknowledgments
This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. RZ, TQ, and SK acknowledge support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. RZ is supported by a Kempner Institute Graduate Research Fellowship, Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. TQ and DAM are partially supported by the Kempner Institute, the Aramont Fellowship Fund, and the FAS Dean’s Competitive Fund for Promising Scholarship. Our work has been improved by invaluable discussion with Will Merrill and David Chiang.
Appendix A Experimental details on synthetic tasks
Below we provide more details when training decoder-only language models on the count and addition—with index hints— tasks. Hyperparameters are largely adapted from [49]. We train all of our models to convergence on the train distribution. Our experiments were run on an internal cluster and all model scales can be run on a single 40GB A100 GPU with gradient accumulation; for count, runs can finish within 2 hours and for addition, runs can finish within 6 hours.
Count task: For all of our training runs, we fix the vocabulary size to 150. For evaluation, we compute the exact match (EM) accuracy across all consecutive subsequences of the test length.
  • Model scales: As mentioned in Section 2, we scale up our models by fixing width and scaling depth and fixing depth and scaling width. The precise parameters for each variation are as follows. For our fixed depth experiments, we fix the network depth to 4 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768,1024\}$. The head dimension is fixed to 64. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,4,6,8\}$ layers.
  • Hyperparameters: We use a learning rate of $31e-$ 3 with a cosine decay scheduler and weight decay 0.1. We set the maximum training duration to be 10000 steps, with batch size 128 and context length 256.
Reverse Order Addition with Index Hints: For evaluation, we compute the exact match (EM) accuracy across 500 batches of 128 examples each.
  • Model scales: For our fixed depth experiments, we fix the network depth to 6 layers and vary width by taking hidden dimensions $\{64,128,256,384,512,640,768\}$. For our fixed width experiments, we fix the hidden dimension to be 512 and vary the depth from $\{1,2,3,4,6,8,10,12\}$ layers.
  • Hyperparameters: We use a learning rate of $41e-$ 4 with a cosine decay scheduler and weight decay 0. We set the maximum training duration to be 30000 steps, with batch size 64 and context length 512.
Appendix B Breakthroughness and Linearity
Srivastava et al. [41] introduced breakthroughness and linearity metrics to capture model performance improving suddenly or reliably with scale. Given a model’s performances $y_{i}$ at model scales $x_{i}$ sorted by ascending model scale, the linearity metric $L$ and breakthroughness metric $B$ are respectively calculated as
$L=\frac{I(y)}{\mathrm{RootMeanSquare}(\{y_{i+1}-y_{i}\}_{i})},$
$B=\frac{I(y)}{\mathrm{RootMedianSquare}(\{y_{i+1}-y_{i}\}_{i})}$
where $I(y)=\mathrm{sign}(\arg\max_{i}y_{i}-\arg\min_{i}y_{i})(\max_{i}y_{i}-\min_{i}y_{i})$.
In Figure 8 we sample the five top seeds for the breakthroughness and linearity metric respectively for count (above) and addition (below).
(a) Count.
(b) Addition.
Figure 8: Top five seeds according to breakthroughness and linearity metrics. Definitions for the two metrics are given in Appendix B, with resulting seeds plotted for 8(a) count and 8(b) addition.
Appendix C Additional Exact Match Accuracy Histogram Plots
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 4 heads per layer.
Figure 9: Random variation in length generalization (count task). Histograms of exact match accuracy on length $60$ sequences when independently scaling 9(a) width and 9(b) depth.
Figure 10: Histograms for EM accuracy when fixing depth to be one layer (count task). We plot the random variation distribution at test length 30 (blue, i.e. in-distribution) and test length 60 (orange, i.e. out-of-distribution). While all model seeds obtain near perfect accuracy in-distribution, all model seeds fail to length generalize at this depth.
Our plots in the main paper fixed test length 40 for addition (Figure 2). In Figure 9 we show analogous histograms for a fixed test length of 60 on the count task. We note that unlike addition, the mode corresponding to failing runs disappears at the highest widths and depths ran, indicating a difference in task difficulty affecting the random variation distribution.
In Section 3.3, we posit that the performance distribution becomes bimodal at the minimum capacity required by the task. For example, Figure 10 shows that, regardless of width, models with a fixed depth of 1 layer are unable to length generalize on the count task, despite achieving near-perfect accuracy in-distribution.
Appendix D Additional Continuous Metric Plots
In Section 3.5 we have seen seen that the addition histograms are bimodal in exact match accuracy (Figure 2) and confirmed that the bimodal random variation distribution persists when viewing a continuous performance metric (Figure 6). We can consider another continuous metric, which corresponds to the minimum probability assigned to a token for each sample in the test length generalization dataset:
$\textrm{minprob}(f,X)=\frac{1}{|X|}\sum_{x\in X}\min_{i<|x|}\mathbf{P}_{f}[x_{i}|x_{0\ldots i-1}]$ (2)
We consider the average minimum probability (and analogously average maximum loss) in the sequence because errors are rare, and thus averaging across the sequence obscures generalization failures. However, as the model improves, the lowest-probability token may shift, but this transition is still continuous, not abrupt. Thus, unlike 0/1 accuracy, this metric also avoids threshold effects and better reflects gradual improvements in length generalization, similar to the continuous error metric. We plot the resulting histograms of this metric in Figures 11 and 12 for addition and count respectively. We note that bimodality persists for both tasks, particularly for addition; in particular, stronger bimodality exhibited in the accuracy histograms corresponds to stronger bimodality in the probability histograms.
In Figure 13, we plot the analogous KDE plot of the loss-based error metric for the count task. For count, we note that the mode corresponding to ‘failing runs’ when looking at accuracy or probability is more diffuse compared to addition. However, we claim that stronger bimodality in the original accuracy/probability plots in Figures 9 and 12 is still associated with stronger bimodality in the loss plots. For the addition task, the scales exhibiting the strongest bimodality in accuracy and probability are for width 4 and 8 (in the case of fixed depth) and depth 2 and 4 (in the case of fixed width), and these are also the most strongly bimodal loss KDE distributions. For the count task, even the most strongly bimodal setting (fixed width, with depth 2) has a low probability of failure ( 25%), but still produces a visible wide peak elevated over the long tail of the loss distribution.
(a) Fixing network depth to 6 layers while scaling up width.
(b) Fixing network width to 8 heads per layer while scaling up depth.
Figure 11: Histograms of the minimum probability of any token in each sequence, averaged across sequences (addition task). Metrics are calculated by Equation 2. The bimodal nature of the random variation distribution persists even when using this continuous metric, analogous to Figure 6.
(a) Fixing network depth to 4 layers.
(b) Fixing network width to 256 hidden dimension.
Figure 12: Histograms of the minimum probability of any token in each sequence, averaged across sequences (count task). Random variation still leads to bimodal performance distributions, even using this continuous performance metric.
(a) Fixing network depth to 4 layers while scaling up width.
(b) Fixing network width to 6 heads per layer while scaling up depth.
Figure 13: Random variation in continuous length generalization error (count task). Kernel Density Estimation (KDE) of loss-based error metric (Equation 1) distribution across model runs. At scales where the EM accuracy distribution is most strongly bimodal, we see that the KDE places the most density at areas of high loss.
Appendix E CSQA LM Results
E.1 Additional Task
In addition to MMLU task, we also reproduce the same data mix experiment on CommonsenseQA (CSQA) [43] task and we include the result in Figure 14. Models yields the same qualitative behaviors as the MMLU task observed in Figure 7.
Figure 14: Violin plots of last-layer reinitialized Qwen2.5 models on the CSQA benchmark. The same behavior observed in MMLU in Figure 7 is reproduced on CSQA as well.
E.2 Impact of Model Size
We repeat the same data mix experiment detailed in Section 2.2 on Qwen2.5-1.5B and report results in Figure 15. In contrast to Figure 7 left, a larger model can learn to perform well on the MMLU task with fewer samples in the training data. For example, Qwen2.5-1.5B can consistently achieve non-trivial performance when the continually pretrained on a data mix that only contains 5% of MMLU auxiliary training samples. In contrast, to achieve a similar performance distribution across 20 runs, Qwen2.5-0.5B requires a data mix that contains 20% of MMLU auxiliary training samples. Interestingly, while the emergence of MMLU capability for Qwen2.5-0.5B showcase bimodal behavior, such a phenomenon is not observed in Qwen2.5-1.5B. As we increase the MMLU ratio in the training data from 0.01% to 0.5%, we instead see a more continuous transition from 0% MMLU performance to non-trivial performance. We qualitatively observe that even with a very small amount of MMLU samples, the last-layer reinitialized Qwen2.5-1.5B model can sometimes learn to output multiple-choice answers but not consistently across all the MMLU test questions. The inconsistent format following behaviors leads to a more continous transition that we observe in Figure 15.
Given that the even the performance distributions with the highest variability in Qwen2.5-1.5B are not clearly bimodal, it appears that bimodality depends on model capacity, at least in our continual pretraining setup. We speculate that even a small number of MMLU examples may draw out a latent ability to handle multiple choice formats in the 1.5B model. There is support for the notion that a model might encode some useful latent structure without yet being able to use it; masked language models learn to represent linguistic structure before they learn to use it for complex grammaticality judgments [3].
Figure 15: Violin plots for last-layer reinitialized Qwen2.5-1.5B models on the MMLU benchmark. The emergence of MMLU capability on Qwen2.5-1.5B shows a more continuous transition.